# On Large Language Models in Mission-Critical IT Governance: Are We Ready Yet?

### v1

@@ Abstract
Context. The security of critical infrastructure has been a fundamental concern since the advent of computers, and this concern has only intensified in today’s cyber warfare landscape. Protecting mission-critical systems (MCSs), including essential assets like healthcare, telecommunications, and military coordination, is vital for national security. These systems require prompt and comprehensive governance to ensure their resilience, yet recent events have shown that meeting these demands is increasingly challenging.
Aim. Building on prior research that demonstrated the potential of GAI, particularly Large Language Models (LLMs), in improving risk analysis tasks, we aim to explore practitioners’ perspectives, specifically developers and security personnel, on using generative AI (GAI) in the governance of IT MCSs seeking to provide insights and recommendations for various stakeholders, including researchers, practitioners, and policymakers.
Method. We designed a survey to collect practical experiences, concerns, and expectations of practitioners who develop and implement security solutions in the context of MCSs. Analyzing this data will help identify key trends, challenges, and opportunities for introducing GAIs in this niche domain.
@@
Mission Critical, IT, Governance, Large Language Model, Security, Survey, XAI.
@@ I Introduction
Since the conception of computers [1, 2], security has been a requirement or an issue [3]. In the current cyber warfare landscape, where wars are fought over fiber optics cable [4], protecting critical infrastructure is becoming more daunting and paramount with each passing second [5]. Currently, our daily activities almost entirely rely on digital platforms [4]. Anyhow, telecommunication infrastructures enabling surfing the internet for a cooking recipe is not as critical or as fault-tolerant as the one used during a national crisis for military personnel and population coordination [6, 7, 8]. We commonly refer to such vital infrastructure as “mission-critical systems” (MCSs), according to the Common Criteria’s definitions [6]. Therefore, the context in which such systems operate is referred to as “mission-critical context” (MCC). MCSs need prompt and actionable plans to safeguard and protect them as soon and as much as possible [3]. Currently, each MCSs follows standardized rules and security practices that vary among nations and according to the criticality level [6, 7, 8] of the system. Hence, the laws and the standards defining the governance of such systems are fragmented and mostly incompatible among the various nations [9].
@@
Nonetheless, recent events exemplified that experts often can not match the three requirements [10, 11]. Stemming from the strict requirements and the latest known failed experiences [10, 11] we previously investigated the role of generative AI (GAI) in aiding human experts performing comprehensive risk analysis, one of the three main tasks of GMCS, according to ANS and ISO standards [3, 10, 12, 13]. Our findings proved that GAI, in the form of Large Language Models (LLMs), can aid, in some instances also, outperform human experts in terms of speed and accuracy.
In recent engagements with the industry, we received exciting opinions from the experts we were collaborating with on the general interest of introducing GAI in GMCS. Hence, to support our future research effort, and contribute to our community, we considered the industrial partner’s initial designs, the ethical implications, and regulations on the topic, and leveraging our experience, we aimed to investigate the practitioner’s point of view on integrating GAIs, specifically LLMs, in the GMCSs. More specifically: (i) we provide the first survey on practitioners feelings and perceptions on integrating GAI in GMCS; (ii) we elaborate implications tailored to researchers to guide future works; (iii) we suggest actionable insights for practitioners; (iv) we provide suggestions for policymakers to craft informed regulations and policies.
@@ II-A Goal and Research Questions
$RQ_{2}$. What are the perceived benefits and potential limitations of LLMs in GMCS?
Research is usually faster than industries when the latter are not involved in research and innovation (RI) initiatives [20]. Therefore, researcher-driven innovation can be far from real-world immediate actionability. Hence, research can lead to exciting results that may be hard to implement in a real-world MCC scenario [21]. In a conservative context such as the GMCS, it is paramount to drive innovation slowly and based on the practitioners’ perception [22]. Therefore, this RQ aims to identify perceptions of benefits and potential limitations in addressing security challenges.
@@
We are in the peculiar spot of a fast-changing regulation due to the impact AI is having on society. Recently, primary regulatory bodies, such as the European Union (EU), published a corpus of regulation, namely the AI Act [27], in which they intended to regulate AI technologies within the perimeter of the EU. It’s one of the first comprehensive efforts by any major regulatory body to address AI’s ethical, legal, and safety implications on a large scale. Conversely, the U.S. approach to AI regulation is more fragmented and decentralized compared to the EU’s comprehensive and unified framework [28, 29, 30, 31]. While the EU AI Act provides a clear, risk-based regulatory structure, the U.S. has so far focused on guidelines, principles, and sector-specific regulations rather than a single, overarching AI law. Therefore, in the context of policies and policymakers, we aim to gather insights from our niche practitioners that can also aid the former in developing or consolidating current frameworks.
@@
RQ1 Q11 How familiar are you with large language models (LLMs) such as ChatGPT (GPT-3/4), BERT et simila? L
@@ II-B Study Context
According to ISACA [32], our work is focused on GMCS, which involves the structured manner of directing, managing, and controlling IT resources to ensure full compliance with organizational objectives, value maximization, and mitigation of risks associated with core operations [9, 3]. Much emphasis is placed on good governance through oversight, accountability, and decision-making to address IT systems’ continued availability, security, and reliability as intrinsic enablers to organizational success. It is that form of governance that is key in ensuring the best performance and compliance with regulatory standards and that the investments being made in technology support the core mission objectives and resilience [9, 30, 10].
@@ II-D Questionnaire
Table I presents the questions induced by our RQs. Due to space constraints, we provide the complete questionnaire in the replication package. It is worth noticing that, before submitting the questionnaire, we have performed a pilot questionnaire to derive the answers for the closed-ended questions (C) predefined answers. Moreover, for most of the closed-ended questions we provided an opened ended version (B) to account for possible missing categories or to allow room for an in-depth explanation of the selection, though we flag them as optional. We interviewed the experts involved in the previous papers [3, 10], and leveraging domain knowledge, we reached a consensus for the predefined answer to the closed-ended questions. We did not ask the pilot’s experts to participate in the final questionnaire to avoid biases.
@@ II-E Data Analysis
Although we already know that our specific context follows mostly a conservative approach towards new technologies, we do not deem that the responses to RQ2, RQ3, and RQ4 may perfectly align with prior expectations, so we anyway adopted an inductive approach to develop a new theory based on the qualitative data provided. The open-ended questions for RQ2 to RQ4 were manually coded in the following as follows. Initially, the first and second authors independently coded the responses for subsets of related questions (RQ2: Q17, RQ3: Q23, and RQ4: Q24, Q25). This process began with open coding, as Strauss and Corbin [35] outlined, followed by axial coding to identify higher-level categories. Subsequently, the two coders compared their coding and discussed any discrepancies until they reached a consensus. The coding process involved assigning codes to small, coherent units within the responses and organizing the emerging concepts (benefits/limitations) into a hierarchical structure of categories. This process was repeated iteratively until a state of saturation was achieved, meaning no new codes or categories were identified. Anyhow, we remark that open-ended questions were optional; thus, at the end of the coding sessions, due to the lack of answers, hence a lack of many categories worth investigating and citing in the results, we deemed to cite the relevant answers as a part of the argumentation in the discussion section where relevant.
II-F Availability and Replicability
We provide all the raw answers, gathered insights, and questionnaires publicly available on our replication package hosted on Zenodo [^1] https://doi.org/10.5281/zenodo.13918422.
@@ III Results
This section presents the results of our survey. We collected data from 31 interviewee. In the following, we first summarize the information about the study population before describing the results for each of the RQs. Due to space constraints, we present the data only in a tabular format, avoiding graphs. We note that for the closed-ended questions we decided to mention only the top-three options.
e Age (Q1) 18-24 6,45% Occupation (Q7) Consultant 6,45% 25-34 12,90% Government Official 38,31% 35-44 32,26% IT/Technology Professional 45,16% 45-54 29,03% Law Enforcement Officer 3,23% 55-64 12,90% Security Analyst 6,45% Over 65 6,45% Industry Sector (Q8) Defense/Security 12,90% Gender (Q2) Female 32,26% Education/Research 3,23% Male 67,74% Government/Public Sector 41,94% Education Level (Q3) Bachelor’s degree 12,90% Healthcare/Pharmaceutical 3,23% Doctoral degree 9,68% Technology/Software 38,71% High school diploma or equivalent 32,26% Years of Experience (Q9) 1-3 12,90% Master’s degree 45,16% 4-6 9,68% Geographical Location (Q4) Europe 83,87% 7-10 9,68% North America 16,13% Over 10 67,74% Organization Size (Q5) Enterprise (10.000+ employees) 12,90% Primary Area of Interest or Expertise (Q10) Artificial Intelligence 3,23% Large (501-10,000 employees) 16,13% Building Design and Planning 3,23% Medium (51-500 employees) 25,81% Cybersecurity 19,35% Small (1-50 employees) 19,35% Data Science/Analytics 9,68% Primary Role in Organization (Q6) Analyst 6,45% Governance/Policy Making 35,48% Executive/Managerial 32,26% Risk Management 9,68% Policy Making/Compliance/Regulatory 32,26% Software Development 3,23% Researcher 3,23% Software Engineering 16,13% Technical Specialist 25,81%
@@ III-B Familiarity and experience with LLMs (RQ1)
Table III shows practitioners’ familiarity with LLMs. According to Table III, most interviewees are slightly to quite familiar with LLMs, with 16% being extremely familiar and $\sim$26having no experience at all with them. Moreover, 32% of the interviewees have specific experiences in employing LLMs for data or risk analysis.
@@
III-C Perceived Benefits and Limitations (RQ2)
Table IV shows the practitioners’ point of view on the future role, the potential limitations, and concerns on LLMs for risk analysis in MCC. According to Table IV, practitioners focused on three main future roles for LLMs in RA (Q15). More spefifically, 27% of the interviewees affirmed that LLMs are useful for automated threat detection and response and enhancing the overall predictive power of RA. Finally, envisioning a continuously running LLM, the interviewee also highlights the potential for real-time anomaly detection. Similarly, practitioners highlighted legal and regulatory compliance challenges, limited contextual understanding, and high computation resource requirements as the top three potential limitations (Q16). Moreover, privacy and data protection, accuracy and reliability of LLMs outputs and potential bias in data analysis where the top three concerns for the practitioners (Q17).
@@ III-C Perceived Benefits and Limitations (RQ2)
Take Aways for RQ2. Practitioners expressed positive perceptions of LLMs in GMCS. However, legal and regulatory compliance challenges and limited contextual understanding are major concerns our community needs to address in the future.
TABLE IV: LLMs in Risk Analysis
@@ III-D Impact on current GMCS frameworks (RQ3)
Table VI presents the impact of integrating LLMs in the currently established governance workflow. According to Table VI, 23% of the interviewees expressed an optimistic view of the integration, and 46% of them stated that LLMs have an innovative potential (Q22). Nonetheless, $\sim$18expressed a cautious approach while 5% a skeptical view (Q22). Furthermore, when asked to agree on the ease of integration (Q23), $\sim$48expressed a neutral agreement, and $\sim$23expressed a good agreement, with 6% strongly agreeing and 6% strongly disagree. Similarly, when asked to agree on the absence of concerns in the integration, (Q24) $\sim$39of the participants had a neutral agreement. In contrast, 19% of them agreed with 6% strongly agreeing and 12% strongly disagree.
@@ III-E Role of policy in LLM’s safe and ethical use (RQ4)
Regarding the necessary regulations for the safe and ethical use of LLMs in MCC governance (Q27), it is worth noticing that, according to the interviewee, there is no clear winner among the options. The top two options are mandating regular audits and compliance checks for LLM deployments ($\sim$29%) and establishing a certification process for LLM technologies (22%). Similarly, regarding methodologies for developing such policies, (Q28), the interviewee suggested forming interdisciplinary committees to draft comprehensive policies ($\sim$30%) and benchmarking policies from other countries already implementing LLMs in their MCC governance (28%).
Regarding data privacy (Q29), the interviewee suggests creating policies enforcing strict data anonymization techniques when LLMs handle sensitive information ($\sim$33%) and developing national and international guidelines for secure data storage and transmission of trained model data (25%).
@@
Futhermore, the interviewee suggested interesting initiatives for Promoting Policy Makers, Practitioners and Researchers’ collaboration on policy-making (Q31), such as hosting regular workshops and seminars with stakeholders ($\sim$29%) and encouraging public-private partnerships for collaborative research (27%).

### v2

@@ Abstract
Context. The security of critical infrastructure has been a pressing concern since the advent of computers and has become even more critical in today’s era of cyber warfare. Protecting mission-critical systems (MCSs), essential for national security, requires swift and robust governance, yet recent events reveal the increasing difficulty of meeting these challenges.
Aim. Building on prior research showcasing the potential of Generative AI (GAI), such as Large Language Models, in enhancing risk analysis, we aim to explore practitioners’ views on integrating GAI into the governance of IT MCSs. Our goal is to provide actionable insights and recommendations for stakeholders, including researchers, practitioners, and policymakers.
Method. We designed a survey to collect practical experiences, concerns, and expectations of practitioners who develop and implement security solutions in the context of MCSs.
@@
Mission Critical, IT, Governance, Large Language Model, Security, Survey, XAI
@@ I Introduction
Since the advent of computers [1, 2], security has been a persistent concern [3]. In today’s era of cyber warfare, where conflicts unfold over fiber optic networks [4], protecting critical infrastructure has become increasingly vital [5]. While daily activities heavily rely on digital platforms [4], telecommunication systems used for casual tasks differ significantly in criticality and fault tolerance from those required for national crises, such as military and population coordination [6, 7, 8]. These vital systems, known as “mission-critical systems” (MCSs)[6], operate within a “mission-critical context” (MCC). MCSs require immediate and robust security measures[3], yet their governance standards remain fragmented and incompatible across nations [9].
@@
Recent events have highlighted that experts often struggle to meet the three core requirements of GMCS [10, 11]. Motivated by these challenges and documented failures [10, 11], we previously explored the potential of Generative AI (GAI) to assist human experts in performing comprehensive risk analysis, one of the key tasks outlined in ANS and ISO standards [3, 10, 12, 13]. Our findings demonstrated that GAI, particularly Large Language Models (LLMs), can support and outperform human experts in speed and accuracy. Therefore, stemming from insights and strong interest from industrial partner regarding the integration of GAI into GMCS [3, 10], we leveraged such insights, preliminary designs, ethics, and regulations on the topic, to drive our contributions to the community in future research. Based on those experiences, we tried to study practitioners’ views about the usage of GAIs, especially LLMs, inside GMCS. More specifically: (i) we provide the first survey on practitioners feelings and perceptions on integrating GAI in GMCS; (ii) we elaborate implications tailored to researchers to guide future works; (iii) we suggest actionable insights for practitioners; (iv) we provide suggestions for policymakers to craft informed regulations and policies.
@@
Data Availability Statement
We publicly provide all raw answers, gathered insights, and questionnaires as part of our replication package, which is hosted on Zenodo [^1] https://doi.org/10.5281/zenodo.13918422.
@@ II-A Goal and Research Questions
$RQ_{2}$. What are the future roles, perceived benefits and potential limitations of LLMs in GMCS?
Research is usually faster than industries when the latter are not involved in research and innovation (RI) initiatives [20]. Therefore, researcher-driven innovation can be far from real-world immediate actionability. Hence, research can lead to exciting results that may be hard to implement in a real-world MCC scenario [21]. In a conservative context such as the GMCS, it is paramount to drive innovation slowly and based on the practitioners’ perception [22]. In our study, we interviewed practitioners who are currently using LLMs as well as those who are not. Therefore, this RQ aims to identify potential future roles of LLMs in GMCS as well perceptions of benefits and potential limitations in addressing security challenges.
@@
We are in the peculiar spot of a fast-changing regulation due to the impact AI is having on society. Recently, primary regulatory bodies, such as the European Union (EU), published a corpus of regulation, namely the AI Act [27], in which they intended to regulate AI technologies within the perimeter of the EU. It’s one of the first comprehensive efforts by any major regulatory body to address AI’s ethical, legal, and safety implications on a large scale. Conversely, the U.S. approach to AI regulation is more fragmented and decentralized compared to the EU’s comprehensive and unified framework [28, 29, 30, 31]. While the EU AI Act provides a clear, risk-based regulatory structure, the U.S. has focused on guidelines, principles, and sector-specific regulations rather than a single, overarching AI law. Therefore, in the context of policies and policymakers, we aim to gather insights that can also aid the former in developing or consolidating current frameworks.
@@
RQ1 Q11 How familiar are you with large language models (LLMs) such as ChatGPT (GPT-3/4), BERT and similar? L
@@ II-B Study Context
Practical Scenario and Previous Insights. This work stems from our previous experience with domain experts tasked to experimentally and gradually introducing LLMs in the context of GMCS [3, 10]. We are in the shoes of an expert tasked with evaluating an IT system in Italy that processes, e.g., a mission critical health information system [3]. The evaluation must align with the regulatory requirements of Regulation (EU) 2016/679 (GDPR) and national laws, addressing both general health data protection and the specifics of electronic health records. Additionally, the expert must accommodate the specific needs of the entity using the system.
The expert’s role involves identifying regulatory and business requirements, gathering relevant documentation, and assessing procedures, staff training, security protocols, physical infrastructure, access controls and the system itself. After synthesizing and analyzing this information, the expert conducts a thorough evaluation, prepares a final report, and outlines any necessary action plans.
Effective Governance. According to ISACA [32], our work is focused on GMCS, which involves the structured manner of directing, managing, and controlling IT resources to ensure full compliance with organizational objectives, value maximization, and mitigation of risks associated with core operations [9, 3]. Much emphasis is placed on good governance through oversight, accountability, and decision-making to address IT systems’ continued availability, security, and reliability as intrinsic enablers to organizational success. It is that form of governance that is key in ensuring the best performance and compliance with regulatory standards and that the investments being made in technology support the core mission objectives and resilience [9, 30, 10].
@@
Motivations for Using LLMs. Unlike financial and credit risk assessment [33, 34, 35], which primarily involves quantitative data, i.e., such as the number of transactions, their sources and destinations, and their content, RA in our context relies heavily on qualitative aspects [10], such as natural language descriptions of scenarios and systems, alongside quantitative data hence the analysis, the anomaly detection and the overall assessment is more challenging. As a result, traditional Machine Learning and other Neural Network models are less suitable for this application compared to LLMs. Despite their limitations, LLMs can significantly enhance data collection and analysis, allowing experts to concentrate on critical tasks like evaluation and action plan development.
@@ II-D Questionnaire
Table I presents the questions induced by our RQs. Due to space constraints, we provide the complete questionnaire in the replication package. It is worth noticing that, before submitting the questionnaire, we have performed a pilot questionnaire to derive the answers for the closed-ended questions (C) predefined answers. Moreover, for most of the closed-ended questions we provided an open-ended version (B) to account for possible missing categories or to allow room for an in-depth explanation of the selection, though we flag them as optional. We interviewed the experts involved in the previous papers [3, 10], and leveraging domain knowledge, we reached a consensus for the predefined answer to the closed-ended questions. We did not ask the pilot’s experts to participate in the final questionnaire to avoid biases.
@@ II-E Data Analysis
Although we already know that our specific context follows mostly a conservative approach towards new technologies, we do not deem that the responses to RQ2, RQ3, and RQ4 may perfectly align with prior expectations, so we anyway adopted an inductive approach to develop a new theory based on the qualitative data provided. The open-ended questions for RQ2 to RQ4 were manually coded as follows. The first and second authors independently coded responses to related questions (RQ2: Q17, RQ3: Q23, and RQ4: Q24, Q25) using open and axial coding, following Strauss and Corbin [38]. We addressed disagreement via discussion, and codes were organized into a hierarchy of benefits and limitations until saturation was reached. Since the open-ended questions were optional and yielded limited responses, we incorporated relevant answers into the discussion section to support the argument where appropriate.
@@ III Results
This section presents the results of our survey. We collected data from 31 interviewees. In the following, we first summarize the information about the study population before describing the results for each of the RQs. Due to space constraints, we present the data only in a tabular format, avoiding graphs. We note that for the closed-ended questions we decided to mention only the top-three options.
@@
Age (Q1) 18-24 6,45% Occupation (Q7) Consultant 6,45%
25-34 12,90% Government Official 38,31%
35-44 32,26% IT/Technology Professional 45,16%
45-54 29,03% Law Enforcement Officer 3,23%
55-64 12,90% Security Analyst 6,45%
Over 65 6,45% Industry Sector (Q8) Defense/Security 12,90%
Gender (Q2) Female 32,26% Education/Research 3,23%
Male 67,74% Government/Public Sector 41,94%
Education Level (Q3) Bachelor’s degree 12,90% Healthcare/Pharmaceutical 3,23%
Doctoral degree 9,68% Technology/Software 38,71%
High school diploma or equivalent 32,26% Years of Experience (Q9) 1-3 12,90%
Master’s degree 45,16% 4-6 9,68%
Geographical Location (Q4) Europe 83,87% 7-10 9,68%
North America 16,13% Over 10 67,74%
Organization Size (Q5) Enterprise (10.000+ employees) 12,90% Primary Area of Interest or Expertise (Q10) Artificial Intelligence 3,23%
Large (501-10,000 employees) 16,13% Building Design and Planning 3,23%
Medium (51-500 employees) 25,81% Cybersecurity 19,35%
Small (1-50 employees) 19,35% Data Science/Analytics 9,68%
Primary Role in Organization (Q6) Analyst 6,45% Governance/Policy Making 35,48%
Executive/Managerial 32,26% Risk Management 9,68%
Policy Making/Compliance/Regulatory 32,26% Software Development 3,23%
Researcher 3,23% Software Engineering 16,13%
Technical Specialist 25,81%
@@ III-B Familiarity and experience with LLMs (RQ1)
Table III shows practitioners’ familiarity with LLMs. According to Table III, most interviewees are slightly to quite familiar with LLMs, with 16% being extremely familiar and $\sim$26having no experience at all with them. This last results was expected because we were interested both in the practitioners that is currently using LLMs and those who aren’t. Moreover, 32% of the interviewees have specific experiences in employing LLMs for data or risk analysis.
@@
III-C Future Roles, Perceived Benefits and Limitations (RQ2)
Table IV highlights practitioners’ perceived benefits, potential limitations, and key concerns surrounding LLMs use in GMCS. According to Table IV, practitioners focused on three main future roles for LLMs in RA (Q15). More specifically, 27% of the interviewees affirmed that LLMs are useful for automated threat detection and response and enhancing the overall predictive power of RA. Finally, envisioning a continuously running LLM, the interviewee also highlights the potential for real-time anomaly detection. Similarly, practitioners highlighted legal and regulatory compliance challenges, limited contextual understanding, and high computation resource requirements as the top three potential limitations (Q16). Moreover, privacy and data protection, accuracy and reliability of LLMs outputs and potential bias in data analysis were the top three concerns for the practitioners (Q17).
@@ III-C Future Roles, Perceived Benefits and Limitations (RQ2)
Take Aways for RQ2. Practitioners expressed positive perceptions on the future roles of LLMs in GMCS. However, legal and regulatory compliance challenges and limited contextual understanding are major concerns our community needs to address in the future.
TABLE IV: LLMs’ Perceived Benefits and Potential Limitations in GMCS
@@ III-D Impact on current GMCS frameworks (RQ3)
Table VI presents the impact of integrating LLMs in the currently established governance workflow. According to Table VI nearly 50% of the interviewees adopted a neutral stance on each question, neither demonizing nor praising LLMs, which reflects their generally conservative approach.
More specifically, 23% of the interviewees expressed an optimistic view of the integration, and 46% of them stated that LLMs have an innovative potential (Q22). Nonetheless, $\sim$18expressed a cautious approach while 5% a skeptical view (Q22). Furthermore, when asked to agree on the ease of integration (Q23), $\sim$48expressed a neutral agreement, and $\sim$23expressed a good agreement, with 6% strongly agreeing and 6% strongly disagree. Similarly, when asked to agree on the absence of concerns in the integration, (Q24) $\sim$39of the participants had a neutral agreement. In contrast, 19% of them agreed with 6% strongly agreeing and 12% strongly disagree.
@@ III-E Role of policy in LLM’s safe and ethical use (RQ4)
Regarding the necessary regulations for the safe and ethical use of LLMs in MCC governance (Q27), it is worth noticing that, according to the interviewees, there is no clear winner among the options. The top two options are mandating regular audits and compliance checks for LLM deployments ($\sim$29%) and establishing a certification process for LLM technologies (22%). Similarly, regarding methodologies for developing such policies, (Q28), the interviewees suggested forming interdisciplinary committees to draft comprehensive policies ($\sim$30%) and benchmarking policies from other countries already implementing LLMs in their MCC governance (28%).
Regarding data privacy (Q29), the interviewees suggests creating policies enforcing strict data anonymization techniques when LLMs handle sensitive information ($\sim$33%) and developing national and international guidelines for secure data storage and transmission of trained model data (25%).
@@
Futhermore, the interviewees suggested interesting initiatives for Promoting Policy Makers, Practitioners and Researchers’ collaboration on policy-making (Q31), such as hosting regular workshops and seminars with stakeholders ($\sim$29%) and encouraging public-private partnerships for collaborative research (27%).